# DNABERT-S: Learning Species-Aware DNA Embedding with Genome Foundation Models

### v1

@@ 5.5 Ablation Study
In this section, we present our ablation studies on DNABERT-S. We perform the ablation study on CAMI2 datasets with both clustering and classification. To validate the effectiveness of curriculum learning, we compare DNABERT-S with three of its variants, each of which is trained purely with the Weight SimCLR [Zhang et al., 2021], i-Mix [Lee et al., 2020], and our proposed Manifold Instance Mixup (MI-Mix) loss. To examine the effectiveness of MI-Mix, we also compare it with a variant trained with the curriculum contrastive method that replaces MI-Mix with i-Mix in the second phase. All the variants are trained with the same data and hyperparameters.
As shown in Table 2, our curriculum learning strategy that combines Weighted SimCLR and MI-Mix achieves the best performance. Our method outperforms both variants that are trained purely with Weighted SimCLR and MI-Mix loss, showing the effectiveness of our proposed curriculum contrastive learning strategy. Moreover, the comparison among the three variants that are trained with a single loss function indicates the effectiveness of MI-Mix in learning DNA embeddings.
@@
W. SimCLR + MI-Mix (ours) $51.11$ $60.88$
W. SimCLR + i-Mix 3.46-3.46 3.56-3.56
only W. SimCLR 1.13-1.13 1.17-1.17
only MI-Mix 0.66-0.66 0.42-0.42
only i-Mix 5.25-5.25 4.76-4.76
Table 2: Ablation study on the Curriculum Contrastive Learning (C2LR) and Manifold Instance Mixup (MI-Mix).
6 Conclusion
We introduced DNABERT-S, a novel genome foundation model designed to generate effective, species-aware DNA embeddings. To facilitate the training of DNABERT-S, we introduce the Manifold Instance Mixup (MI-Mix) training objective and the Curriculum Contrastive Learning (C2LR) strategy. We perform extensive experiments on $18$ datasets across a variety of challenging tasks, including species clustering, classification, and metagenomics binning, to demonstrate the remarkable ability of DNABERT-S.
Limitations and Broader Impact. We expect the DNABERT-S to greatly benefit a wide range of species-related research, including but not limited to species identification, metagenomics binning, biodiversity assessment, and understanding of evolutionary relationships. Despite its specialized focus on species-aware tasks, as noted in Appendix C.6, DNABERT-S’s species-aware training approach does not inherently enhance its performance in unrelated genomic tasks, such as promoter prediction in the human genome. Nevertheless, the methodologies and insights gleaned from our extensive empirical analyses offer valuable pathways for advancing more detailed and accurate investigations across a wide spectrum of genomic applications.
Acknowledgments
Appendix A i-Mix for Original Contrastive Learning
The Weighted SimCLR method treat each instance $\{x_{i}\}_{i=1}^{B}$ and $\{x_{i^{+}}\}_{i=1}^{B}$ as anchors, with each anchor associated with 2B22B-2 negative samples. This section explains why integrating the i-Mix method into Weighted SimCLR (Sec. 3.1) doubles memory or training time.
[…]
Appendix B Data Statistics of Evaluation Benchmark
Tasks Source ID Species Sequences Sequence Length Sequences Per Species
Synthetic 0 200 20000 10k / 10k / 10k 100 / 100 / 100
Synthetic 1 200 20000 10k / 10k / 10k 100 / 100 / 100
Marine 0 326 32600 2k / 20k / 7.6k 100 / 100 / 100
Marine 1 375 37500 2k / 20k / 8.2k 100 / 100 / 100
Clustering Marine 2 361 36100 2k / 20k / 8.5k 100 / 100 / 100
& Marine 3 499 49900 2k / 20k / 6.8k 100 / 100 / 100
Marine 4 360 36000 2k / 20k / 7.1k 100 / 100 / 100
Classification Plant 0 108 10800 2k / 20k / 6.6k 100 / 100 / 100
Plant 1 100 10000 2k / 20k / 6.4k 100 / 100 / 100
Plant 2 93 9300 2k / 20k / 6.2k 100 / 100 / 100
Plant 3 129 12900 2k / 20k / 5.5k 100 / 100 / 100
Plant 4 129 12900 2k / 20k / 5.7k 100 / 100 / 100
Synthetic 5 323 37278 10k / 10k / 10k 31 / 200 / 111
Synthetic 6 249 28206 10k / 10k / 10k 30 / 199 / 114
Metagenomics Marine 5 515 119465 2.5k / 20k / 4.3k 10 / 841 / 201
Binning Marine 6 527 125194 2.5k / 20k / 4.4k 10 / 915 / 223
Plant 5 181 71642 2.5k / 20k / 3.7k 10 / 4293 / 190
Plant 6 196 68426 2.5k / 20k / 3.7k 10 / 4599 / 116
Appendix C More Experimental Results
C.1 Comparison Between Different Foundation Model
@@
This subsection delineates a comparative analysis of various existing genome foundation models in the context of DNA embedding generation. We evaluated four renowned models: DNABERT [Ji et al., 2021], DNABERT-2 [Zhou et al., 2023], Nucleotide Transformer [Dalla-Torre et al., 2023], and HyenaDNA [Nguyen et al., 2023]. Notably, DNABERT and the Nucleotide Transformer exhibit strict input sequence length limitations of 512 and 6144 (V1) or 12288 (V2) base pairs, respectively. Conversely, DNABERT-2 and HyenaDNA do not impose such constraints. Considering the potentially extensive length of genome sequences in metagenomics binning, our preliminary experiments focused solely on DNABERT-2 and HyenaDNA.
@@
We train both models on our pre-training datasets with the same set of hyperparameters for $3$ epoch. We save checkpoints periodically and select the best checkpoint based on the models’ validation loss on the validation set. Since HyenaDNA, in general, requires larger learning rates than DNABERT-2, we train it with three different learning rates (3e43e-4, 3e53e-5, and 3e63e-6) and select the one that works best. For DNABERT-2, we only train it once with a learning rate of 3e63e-6. To avoid the impact of other factors such as the schedule of curriculum learning, we train both models with the Weighted SimCLR loss only in the entire training process. We then evaluate the models before and after contrastive training on our evaluation benchmark.
@@
Table 4 presents the performance of both pre-trained DNABERT-2 and HyenaDNA, with and without contrastive training, in K-means clustering evaluations. DNABERT-2, despite underperforming without contrastive training, demonstrated superior performance compared to HyenaDNA post-training, highlighting its efficacy in learning effective DNA embeddings. Similar trends are observed in few-shot species classification scenarios, as detailed in Table 5. While DNABERT-2 initially exhibited subpar performance, it significantly improved post-contrastive training, achieving the best results. Thus, DNABERT-2 was chosen as the foundational model for DNABERT-S.
@@
Table 4: Performance of DNABERT-2 and HyenaDNA on K-Means clustering measured by Adjusted Rand Index (ARI). Here, $\Delta$ represents the model’s performance improvement after contrastive training,
@@
Table 5: Performance of DNABERT-2 and HyenaDNA on few-shot classification measured by Macro F1 score. Here, $\Delta$ represents the model’s performance improvement after contrastive training,
@@
C.2 Remaining Results on Species Classification
@@
Figure 5: Results of species classification on other 66 datasets.
C.3 Impact of Sequence Length

### v3

@@ 5.5 Ablation Study
In this section, we present our ablation studies on DNABERT-S. We perform the ablation study on CAMI2 datasets with both clustering and classification. To validate the effectiveness of curriculum learning, we compare DNABERT-S with three of its variants, each of which is trained purely with the Weight SimCLR [Zhang et al., 2021], i-Mix [Lee et al., 2020], SupCon [Khosla et al., 2021], and our proposed Manifold Instance Mixup (MI-Mix) loss. To examine the effectiveness of MI-Mix, we also compare it with a variant trained with the curriculum contrastive method that replaces MI-Mix with i-Mix in the second phase. All the variants are trained with the same data and hyperparameters.
Table 2: Ablation study on the Curriculum Contrastive Learning (C2LR) and Manifold Instance Mixup (MI-Mix). -: Performance difference to W. SimCLR + MI-Mix.
@@
W. SimCLR + MI-Mix $51.11$ $60.88$
W. SimCLR + i-Mix $-3.46$ $-3.56$
only W. SimCLR $-1.13$ $-1.17$
only MI-Mix $-0.66$ $-0.42$
only i-Mix $-5.25$ $-4.76$
only SupCon $-6.75$ $-3.83$
As shown in Table 2, our curriculum learning strategy that combines Weighted SimCLR and MI-Mix achieves the best performance. Our method outperforms both variants that are trained purely with Weighted SimCLR and MI-Mix loss, showing the effectiveness of our proposed curriculum contrastive learning strategy. Moreover, the comparison among the four variants that are trained with a single loss function, including Weight SimCLR, i-Mix, and SupCon, indicates the effectiveness of MI-Mix in learning DNA embeddings.
5.6 Selection of Backbone Model
@@
This subsection delineates a comparative analysis of existing genome foundation models in the context of DNA embedding generation. We evaluated four renowned models: DNABERT [Ji et al., 2021], DNABERT-2 [Zhou et al., 2023], Nucleotide Transformer [Dalla-Torre et al., 2023], and HyenaDNA [Nguyen et al., 2023]. Notably, DNABERT and the Nucleotide Transformer exhibit strict input sequence length limitations of 512 and 6144 (V1) or 12288 (V2) base pairs, respectively. Conversely, DNABERT-2 and HyenaDNA do not impose such constraints. Considering the potentially extensive length of genome sequences in metagenomics binning, our preliminary experiments focused solely on DNABERT-2 and HyenaDNA.
@@
We train both models on our pre-training datasets with the same set of hyperparameters for $3$ epoch. We save checkpoints periodically and select the best checkpoint based on the models’ validation loss on the validation set. Since HyenaDNA, in general, requires larger learning rates than DNABERT-2, we train it with three different learning rates ($43e-$ 4, $53e-$ 5, and $63e-$ 6) and select the one that works best. For DNABERT-2, we only train it once with a learning rate of $63e-$ 6. To avoid the impact of other factors, such as the schedule of curriculum learning, we train both models with the Weighted SimCLR loss only in the entire training process. We evaluate the models before and after contrastive training on our evaluation benchmark.
@@
Table 3 presents the performance of both pre-trained DNABERT-2 and HyenaDNA, with and without contrastive training, in K-means clustering evaluations. DNABERT-2, despite underperforming without contrastive training, demonstrated superior performance compared to HyenaDNA post-training, highlighting its efficacy in learning effective DNA embeddings. Similar trends are observed in few-shot species classification scenarios, as detailed in Table 4. While DNABERT-2 initially exhibited subpar performance, it largely improved post-contrastive training, achieving the best results. Thus, DNABERT-2 was chosen as the backbone model for DNABERT-S.
@@
Table 3: Performance of DNABERT-2 and HyenaDNA on K-Means clustering measured by Adjusted Rand Index (ARI). $\Delta$: the model’s performance improvement after contrastive training,
@@
Table 4: Performance of DNABERT-2 and HyenaDNA on few-shot classification measured by Macro F1 score. $\Delta$: the model’s performance improvement after contrastive training,
@@
6 Conclusion
We introduce DNABERT-S, a model tailored for species-aware DNA embeddings, which is empowered by the proposed Manifold Instance Mixup (MI-Mix) training objective and the Curriculum Contrastive Learning (C2LR) strategy. We perform extensive experiments on $23$ datasets across thousands of different species and a variety of challenging tasks, including species clustering, classification, and metagenomics binning, to demonstrate the DNABERT-S’s capability of species differentiation. Furthermore, we conduct a series of experiments on the training objective, backbone model selection, impacts of sequence length, and feature dimension to provide empirical insights for DNA embedding learning. We also compare DNABERT-S with traditional technique like MMSeq2 in the problem of species differentiation and demonstrate DNABERT-S can achieves slightly better performance with less amount of labeled samples. We envision DNABERT-S to potentially change the way genomic problems are approached from an embedding perspective. The primary limitation of DNABERT-S lies in its high computational demands, a common trait among deep learning models, when compared to more traditional, lightweight methods like Tetranucleotide Frequencies (TNF).
Acknowledgments
Appendix A Algorithm for Metagenomics Binning
Algorithm 1 describes the unsupervised clustering algorithm we used for metagenomics binning, where $\texttt{s}({E_{i},E_{j}})$ represents the cosine similarity of two vectors $E_{i}$ and $E_{j}$. Selection of threshold $\gamma$. As shown in Algorithm 1, the threshold $\gamma$ is the most important hyperparameter that greatly impacts the final binning results. A high threshold results in small and dense clusters while a low threshold results in large yet sparse clusters. Since different models generate embeddings with distinct distributions, a fixed threshold (e.g., $0.9$) could be too high for one model yet too low for another one. In practice, massive hyperparameter searches are needed to determine the best threshold for each model on different datasets. Due to the large size of our experiments and the various types of models we used, an automatic way is needed to fairly choose the threshold for each model on each dataset. For each metagenomics binning dataset, we use the dataset from the same source (e.g., Marine) as it with ID 00 to compute a threshold for each model on it, Specifically, we generate embeddings for each DNA sequence in the dataset and compute the similarities between each DNA sequence and its species center (i.e., the average of all the DNA sequence belongs to this species). The $70$ percentile of all the similarities is used as the threshold. Other hyperparameters. We set minimum bin size $m=10$, number of steps $Z=1000$, and number of iterations $T=3$. We also experimented with $T=3,4,5$ and $60,70,80,90$ percentile of all the similarities is used as the threshold $\gamma$, and found that the results are robust to these hyperparameters.
Algorithm 1 Modified $K$-Medoid Clustering
2:  Initialize: predictions $\bm{p}\in\mathbb{R}^{N},p_{i}=-1\text{ for }i=1,\ldots,N$, similarity matrix $\bm{S}=\bm{E}\bm{E}^{\top}$ with $S_{ij}=0$ if $S_{ij}<\gamma$, density vector $\bm{d}\in\mathbb{R}^{N}$ with $d_{i}=\sum_{j=1}^{N}S_{ij}$
3:  for step $z=1$ to $Z$ do
4:     Select seed index $s=\arg\max_{s\textquoteright}d_{s\textquoteright}$ and corresponding seed $E_{s}$
5:     for iteration $t=1$ to $T$ do
7:        Update seed: $E_{s}\leftarrow\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}E_{i}$
8:     end for
9:     Set $p_{i}\leftarrow z,d_{i}\leftarrow 0$ for each $i\in\mathcal{I}$
10:     Set $d_{x}\leftarrow d_{x}-\sum_{i\in\mathcal{I}}S_{xi}$ for each $x\in[1,2,\ldots,N]$
11:  end for
12:  for step $z=1$ to $Z$ do
13:     Find indices $\mathcal{I}$ where $p_{i}=z$ for each $i\in\mathcal{I}$
14:     if $|\mathcal{I}|<m$ then
15:        Set $p_{i}\leftarrow-1$ for each $i\in\mathcal{I}$
16:     end if
17:  end forReturn: predictions $\bm{p}$
Appendix B Data Statistics of Evaluation Benchmark
Our benchmark contains assets from GenBank [Benson et al., 2012] (license: Creative Commons Attribution Non-Commercial License http://creativecommons.org/licenses/by-nc/2.0/uk/) and CAMI2 [Meyer et al., 2022] (license: Creative Commons Attribution 4.0 International License http://creativecommons.org/licenses/by/4.0/). It is worth noting that both GenBank and CAMI2 preprocess the RNA sequences into DNA equivalents by replacing U with T. Thus, although we did not explicitly analyze RNA sequences, many RNA viruses are considered in both model training and evaluation.
Tasks Source ID Species Sequences Sequence Length Num. Per Species
Marine 0 326 32600 2k / 20k / 7.6k 100 / 100 / 100
Unsupervised Marine 1 375 37500 2k / 20k / 8.2k 100 / 100 / 100
Clustering Marine 2 361 36100 2k / 20k / 8.5k 100 / 100 / 100
& Marine 3 499 49900 2k / 20k / 6.8k 100 / 100 / 100
Few-Shot Marine 4 360 36000 2k / 20k / 7.1k 100 / 100 / 100
Classification Plant 0 108 10800 2k / 20k / 6.6k 100 / 100 / 100
Plant 1 100 10000 2k / 20k / 6.4k 100 / 100 / 100
Plant 2 93 9300 2k / 20k / 6.2k 100 / 100 / 100
Plant 3 129 12900 2k / 20k / 5.5k 100 / 100 / 100
Plant 4 129 12900 2k / 20k / 5.7k 100 / 100 / 100
(Microbe) Synthetic 0 200 20000 10k / 10k / 10k 100 / 100 / 100
Synthetic 1 200 20000 10k / 10k / 10k 100 / 100 / 100
(Mammalian, Synthetic 2 210 21000 10k / 10k / 10k 100 / 100 / 100
Invertebrate, Synthetic 3 210 21000 10k / 10k / 10k 100 / 100 / 100
Protozoa) Synthetic 4 210 21000 10k / 10k / 10k 100 / 100 / 100
Synthetic 5 323 37278 10k / 10k / 10k 31 / 200 / 111
Synthetic 6 249 28206 10k / 10k / 10k 30 / 199 / 114
Metagenomics Marine 5 515 119465 2.5k / 20k / 4.3k 10 / 841 / 201
Binning Marine 6 527 125194 2.5k / 20k / 4.4k 10 / 915 / 223
(Microbe) Plant 5 181 71642 2.5k / 20k / 3.7k 10 / 4293 / 190
Plant 6 196 68426 2.5k / 20k / 3.7k 10 / 4599 / 116
Classification Synthetic 7 200 95309 10k / 10k / 10k 452 / 600 / 600
(Microbe) Synthetic 8 200 95475 10k / 10k / 10k 442 / 600 / 600
Appendix C More Experimental Results
C.1 Results Comparison with Linear and Non-Linear Classifiers
Table 6: DNABERT-S’s performance of species classification with varying numbers of training samples on $3$ datasets: Beyond using logistic regression (LR), we also train a Multi-Layer Perceptron (MLP) with non-linear activation function (ReLU). The term "Difference" denotes the performance gap between "DNABERT-S" and the "best-baseline". The results show that DNABERT-S embedding outperforms the best-existing baseline consistently in both linear and non-linear discriminativity.
Marine 0 Plant 0 Dataset ID 1 2 5 10 20 1 2 5 10 20 LR: best-baseline 27.65 38.81 52.4 58.86 63.29 28.15 36.97 48.2 55.24 60.04 LR: DNABERT-S 50.25 59.41 66.07 68.92 70.75 47.83 55.83 63.01 67.12 69.82 LR: Difference 22.60 20.60 13.67 10.06 7.46 19.68 18.86 14.81 11.88 9.78 MLP: best-baseline 26.07 38.59 53.86 60.27 63.07 27.39 36.19 47.5 54.03 59.13 MLP: DNABERT-S 48.55 59.25 66.09 68.95 69.99 45.31 55.34 63.25 67.25 70.00 MLP: Difference 22.48 20.66 12.23 8.68 6.92 17.92 19.15 15.75 13.22 10.87
Table 7: DNABERT-S’s performance of species classification with varying numbers of training samples on $3$ datasets: Beyond using logistic regression (LR), we also train a Multi-Layer Perceptron (MLP) with non-linear activation function (ReLU). The term "Difference" denotes the performance gap between "DNABERT-S" and the "best-baseline". The results show that DNABERT-S embedding outperforms the best-existing baseline consistently in both linear and non-linear discriminativity.
Synthetic 0 Dataset ID 1 2 5 10 20 LR: best-baseline 44.07 56.11 68.69 75.34 79.54 LR: DNABERT-S 71.36 77.93 83.37 85.81 87.77 LR: Difference 27.29 21.82 14.68 10.47 8.23 MLP: best-baseline 40.28 54.95 69.59 76.54 81.33 MLP: DNABERT-S 68.96 77.47 83.44 85.86 87.77 MLP: Difference 28.68 22.52 13.85 9.32 6.44
C.2 Results with Error Bars
In this section, we present detailed results on species clustering and few-shot classification on DNABERT-S and the most competitive baseline models. Table 8, 9, and 10 respectively show the models’ mean and std on each setting across 5 random seeds. As shown in the tables, DNABERT-S consistently outperforms the baselines with small variances.
Table 8: Performance of models with error bars on Synthetic:0 dataset. We evaluate models using K-Means clustering (Sec. 5.2) and 1/2/5/10/20-shot classification (Sec. 5.4).
Synthetic 0
ARI 1 2 5 10 20
TNF 38.18 ± 1.27 44.30 ± 0.97 56.13 ± 1.10 68.68 ± 0.73 75.24 ± 0.37 79.48 ± 0.07
TNF-K 36.11 ± 0.09 39.51 ± 0.35 50.26 ± 0.89 62.43 ± 0.41 68.53 ± 0.50 72.95 ± 0.36
HyenaDNA 20.10 ± 0.50 30.03 ± 0.49 41.21 ± 0.92 54.42 ± 0.73 63.79 ± 0.59 70.53 ± 0.25
DNA-Double 34.91 ± 0.90 34.61 ± 0.91 46.79 ± 0.39 59.92 ± 0.53 67.45 ± 0.20 73.64 ± 0.25
DNABERT-S 66.94 ± 2.07 71.54 ± 0.51 $7$ 0.68 83.12 ± 0.18 85.63 ± 0.18 87.68 ± 0.15
Table 9: Performance of models with error bars on Marine:0 dataset. We evaluate models using K-Means clustering (Sec. 5.2) and 1/2/5/10/20-shot classification (Sec. 5.4).
Marine 0
ARI 1 2 5 10 20
TNF 24.78 ± 0.23 27.89 ± 0.95 38.69 ± 0.04 52.36 ± 0.26 58.95 ± 0.05 62.69 ± 0.10
TNF-K 25.44 ± 0.50 22.82 ± 0.35 30.06 ± 0.61 40.50 ± 0.14 45.20 ± 0.25 49.35 ± 0.15
HyenaDNA 16.31 ± 0.07 23.89 ± 0.83 33.59 ± 0.01 47.50 ± 0.39 55.61 ± 0.18 61.75 ± 0.13
DNA-Double 26.82 ± 0.45 26.87 ± 1.16 36.98 ± 0.21 49.86 ± 0.12 57.93 ± 0.26 63.33 ± 0.06
DNABERT-S 53.91 ± 0.22 50.37 ± 0.74 59.71 ± 0.11 66.03 ± 0.11 69.00 ± 0.19 70.75 ± 0.13
Table 10: Performance of models with error bars on Plant:0 dataset. We evaluate models using K-Means clustering (Sec. 5.2) and 1/2/5/10/20-shot classification (Sec. 5.4).
Plant 0
ARI 1 2 5 10 20
TNF 26.10 ± 0.70 24.07 ± 0.40 32.04 ± 1.00 43.35 ± 0.24 48.92 ± 0.27 53.29 ± 0.39
TNF-K 25.83 ± 0.67 22.82 ± 0.35 30.06 ± 0.61 40.50 ± 0.14 45.20 ± 0.25 49.35 ± 0.15
HyenaDNA 24.61 ± 0.77 28.45 ± 0.30 36.46 ± 0.61 48.56 ± 0.30 55.13 ± 0.53 59.80 ± 0.27
DNA-Double 22.10 ± 0.46 24.80 ± 0.56 32.91 ± 0.52 44.82 ± 0.18 52.77 ± 0.48 59.35 ± 0.37
DNABERT-S 51.15 ± 1.13 48.39 ± 1.33 55.92 ± 0.83 62.97 ± 0.36 67.14 ± 0.41 69.64 ± 0.15
C.3 Remaining Results on Species Classification
@@
Figure 5: Results of species classification on other $6$ datasets.
C.4 Comparison with alignment-based method
While previous experiments have highlighted DNABERT-S’s exceptional performance in scenarios with limited or no labeled data, this section focuses on its effectiveness in situations where abundant labeled data is available. Specifically, we aim to understand the embedding-based species differentiation method in scenarios where reference genomes of the species to classify are available. We compare embedding-based methods with MMseqs2 [Steinegger and Söding, 2017], a leading alignment-based species classification tool.
For a fair comparison with MMseqs2, which relies on the reference genomes of each species when performing classification, we construct two datasets, each consisting of 200 distinct species. To mimic real-world setups, instead of classifying segment of reference genomes, we simulate 600 long-reads with PBSIM2 [Ono et al., 2021] from each selected species, 100 as the test set and 500 as the training set. For the embedding-based methods, such as DNABERT-S and TNF, we generate embedding for all these sequences and use K-Neareast-Neighbor (KNN) classifier with n equals to 5 for species classification. We respectively use 100, 200, 300, 400, and 500 sequences from each species to construct the training set. For the MMseqs2, we respectively use the reference genome and the simulated long-read sequences as the reference for alignment-based classification.
(a) Synthetic 7
(b) Synthetic 7
Figure 6: Results on species classification when reference genomes are available.
Figure 6 the results of the models on the datasets. As shown in the figure, DNABERT-S starts to outperforms MMseqs2 with 200-300 labeled sequence from each species with a simple KNN classier, while TNF achieves comparable performance as MMseqs2 with about 500 labels sequence per species. These results indicate the potential of embedding-based methods to replace traditional alignment-based method in species classification in data abundant scenarios. Yet a fact that cannot be ignored is that DNABERT-S is much more computational cost than MMseqs2. As a comparison, MMseqs2 requires about 30 seconds on a single CPU to make predictions while DNABERT-S requires about 1 hour on 2 NVIDIA A100 GPUs to do the same thing. Therefore, there is a long way to go to fully replace alignment-based methods in high-throughput genomics analysis.
C.5 Results on Non-Microbe Species
Since DNABERT-S is trained on microbe species (e.g., viruses, fungi, and bacteria), a natural question is whether it can distinguish genomics sequences from species that are largely different from the ones in the training set. To answer this question, we construct 3 synthetic datasets (ID: 2, 3, and 4) that include genomes from invertebrate, protozoa, and mammalian species. We randomly select 70 species from each category to achieve 210 species in total. We perform the same clustering and few-shot classification as presented in Table 1 and Figure 4.
Table 11: Performance of models on non-microbe species. We evaluate models on each dataset using K-Means clustering (Sec. 5.2) and 1/5/20-shot classification (Sec. 5.4).
Synthetic:2 Synthetic:3 Synthetic:4
Dataset ID ARI 1 5 20 ARI 1 5 20 ARI 1 5 20
TNF 20.70 26.28 45.71 56.24 19.11 25.05 42.69 52.27 21.49 25.94 44.96 54.60
TNF-K 18.63 22.42 40.07 50.75 16.91 20.38 36.48 46.74 19.30 21.38 39.23 48.94
HyenaDNA 11.20 17.29 34.77 48.76 10.86 16.50 32.69 45.26 11.58 17.30 35.38 47.86
DNABERT-2 9.18 15.91 33.75 49.00 8.99 15.28 31.03 45.17 9.79 15.38 33.75 47.78
DNA-Double 13.01 16.04 30.56 42.11 12.80 15.91 28.82 39.28 14.15 16.20 31.22 42.25
DNABERT-S 32.70 33.21 49.78 59.01 29.44 29.78 45.65 54.63 33.60 32.58 49.67 57.97
As shown in Table 11, DNABERT-S consistently outperforms baselines across all the datasets and evaluation scenarios, indicating DNABERT-S’s transferability and robustness on species that are significantly different from its training set. However, the improvements over the baselines are less significant, and the absolute scores, such as ARI in clustering and F1 in classification, are also lower than those in microbe datasets. On the one hand, there are higher genetic similarities among mammals compared to the often more significant genetic diversity found in microbes, making it more challenging to distinguish different mammalian species. On the other hand, due to the significant distinction between microbe and mammalian/protozoa species, some of the differentiation rules and markers learned from microbe genomes may not be applicable to genomes of species in other categories, which also suggests the needs of in-domain species-aware training.
C.6 Impact of Sequence Length